x+2y-4=0 i don't understand how to solve it ?!
Help & thanks

Answer by rothauserc(4718)

the standard formula for a line is y = mx +b where m is the slope and b is the y intercept
we are given
x +2y -4 = 0
2y = -x +4
y = -x/2 +2